P1= $8/per footlong,

Qd1= 20

P2= $5 / per footlong

Qd2= 40

Interpret the results. Is this product has elastic or inelastic price elasticity of demand? If you are the seller of this product, which unit price will sell it for to maximize your sells revenue?

Answer in words please
Answer #1

Answer : Here the product has elastic price elasticity of demand. Because here when price falls from $8 to $5 then the quantity demanded increase by more than the price changed level. This means that the demand for the product is elastic.

In case of $8 price level, revenue = price * quantity = 8 * 20 = $160.

In case of $5 price level, revenue = 5 * 40 = $200

As at $5 price level the revenue is highest hence $5 will maximize the revenue.